Problem 1 A researcher sends a questionnaire to 9,000 individuals. She gets back 6,000 responses. The percentage of women among the respondents is 20%. Because of nonresponse, the fraction of women among recipients of the survey could be anything between (low) (Q1) 0.133333 and (high) (Q2)

Please solve the answer for Q2. (The answer is not 0.8)
Answer #1

The answer is 4200/9000 = 0.4666667.

The percentage of women among the 6000 respondents = 20

Thus, number of women respondents = 0.2 × 6000 = 1200.

Now, among the remaining 3000 non respondent individuals, a minimum of 0 and maximum of 3000 can be woman.

Thus, maximum number of women among recipients can be 1200 + 3000 = 4200. And minimum number can be 1200 + 0 = 1200

Thus the fraction of women among recipients is

Q1 = 1200/9000 = 0.133333

and Q2 = 4200/9000 = 0.4666667